How much heat is required to raise the temperature of 400g of ethyl alcohol from 25 Coto 50 Co? The specific heat of ethyl alcohol is 0.58 cal/goC.
Answer #1

Given : m= 400 g , cp = 0.58 cal/g°c , \Delta T = 50°c - 25°c = 25°c

Solution:

Amount of heat required to raise the temperature is given by :

Q = mcp\DeltaT

= (400 g)(0.58 cal/g°c)(25°c)

= 5800 cal

We know that , 1 cal = 4.184 J

So Q = 5800*4.184= 24267.2 J = 24.27 kJ

Hence amount of heat required is : 5800 cal or 24.27 kJ